# Pinching Antenna Systems (PASS): Architecture Designs, Opportunities, and Outlook

### v1

@@ IV-A PASS for Ubiquitous Indoor and Outdoor Coverage
The simple structure of PASS enables it to be easily deployed on diverse structures in both outdoor and indoor scenarios, such as building facades, rooftops, roadside guardrails, and indoor ceilings or walls, as depicted in Fig. 4. In contrast to conventional wireless communications that rely on remotely distributed BSs, where communication links are often partially or totally blocked by obstacles and suffer from significant long-distance path loss, PASS establishes LoS-dominated and short-distance wireless channels with the widespread deployment of waveguides and the optimal activation of PAs near the terminals. By preventing wireless signals from being wasted or even generating interference to others, PASS is capable of facilitating ubiquitous outdoor and indoor coverage in a cost-efficient and energy-efficient manner. Several attractive applications of PASS in indoor and outdoor scenarios will be discussed below.
On the one hand, PASS is quite appealing in indoor scenarios. As shown in the top right of Fig. 4, PASS installed on the ceiling can optimally activate PAs to provide wireless communications for users in different rooms. This mitigates the severe penetration losses that wireless signals from outdoor BSs or indoor access points (APs) experience, particularly at mmWave and THz frequencies. Moreover, by exploiting the near-field channel between PAs and users, efficient beam focusing can be achieved. This not only enhances the desired signal strength but also reduces interference to other indoor users. On the other hand, the widely distributed PASS can be utilized in outdoor mobile scenarios, such as aerial/space communications and vehicular communications, as illustrated in the middle of Fig. 4. For example, recall the fact that conventional BSs with down-tilt antennas are typically inefficient for aerial communications. As a remedy, PASS installed on rooftops and facades can realize tailored upward beamforming for communications and sensing to effectively support medium- and high-altitude aircrafts as well as low-altitude unmanned aerial vehicles (UAVs) or other high-altitude platforms. As a further advancement, PASS can be employed in space communications. By integrating multiple waveguides and numerous PAs, PASS can play as a large-scale antenna array to combat the severe path loss typically encountered in high-frequency satellite communications. Moreover, with PASS uniformly deployed along roadside guardrails, efficient vehicular communications can be achieved by adaptively activating PAs in response to vehicular movement. This is particularly beneficial for autonomous driving in future wireless networks.
@@ IV-B Interplay between PASS and NOMA
The combination of PASS and non-orthogonal multiple access (NOMA) provides a ‘win-win’ solution for 6G and beyond. On the one hand, since multiple PAs on a specific waveguide can only be fed with the same set of signals, how to serve multiple users simultaneously becomes a critical challenge. To address this, it is natural to consider using NOMA to multiplex multiple users in the power domain over the same radio resource [9]. On the other hand, PASS also facilitates flexible NOMA operations. Recall the fact that NOMA relies on successive interference cancellation (SIC) to handle co-channel interference and therefore SIC order significantly affects the communication performance achieved by each NOMA user. Conventionally, a fixed distance-based SIC order is employed for NOMA, where users closer to the BS are assigned higher SIC orders due to better channel conditions. In contrast, as shown in the bottom right of Fig. 4, PASS-based NOMA enables a more flexible SIC operation. This is achieved by dynamically adjusting users’ channel conditions through the allocation of PAs and the design of their activation locations. For example, users originally farther from the BS/AP can have higher channel conditions with more numbers of PAs allocated. By doing so, more customized wireless services can be provided by the PASS-based NOMA.
@@ IV-C PASS for NISE/ISAC
In addition to enhancing communications, PASS can significantly improve the quality and coverage of sensing. Since the links between PAs and sensing targets follow the near-field model, PASS enables near-field sensing (NISE) to precisely estimate angular and range information, as well as the velocity of sensing targets [15]. As shown in the top left of Fig. 4, two structures can be employed in practice, namely the ISAC-PASS structure and the NISE-PASS structure. (1) By allocating part of the waveguides to transmit communication signals and the rest to receive echo-sensing signals, ISAC-PASS can simultaneously provide near-field communication and sensing services within the area of interest. (2) Equipped exclusively with receiving waveguides, NISE-PASS can operate in a low-power mode to facilitate NISE. By leveraging the widespread deployment of PASS and successively/periodically activating sensing PAs, ubiquitous and accurate NISE can be achieved. This is particularly beneficial for practical applications, such as tracking target trajectories or monitoring malicious targets.
@@ IV-D PASS for WPT
Wireless power transfer (WPT) is an efficient technology that uses RF signals to supply power to energy-constrained devices and extend their operational lifetime, especially for Internet-of-Things (IoT) and sensor networks. However, conventional WPT faces the challenge of ‘near-far unfairness’, where users located farther from the wireless power-supply BS/AP experience lower energy-harvesting efficiency due to the degraded channel condition. As shown in the middle right of Fig. 4, this issue can be resolved through PASS-empowered WPT. In particular, tailored PAs can be activated right to each device, which ensures high-quality WPT for all devices, regardless of their locations.
@@ IV-E PASS for PLS
PASS can also enhance secure information transmission in the context of physical layer security (PLS). As shown in the bottom left of Fig. 4, PASS enables legitimate information transmission to effectively ‘bypass’ the eavesdropper. On the one hand, since signal transmission predominantly occurs within the waveguide of PASS, the likelihood of legitimate information being eavesdropped is significantly reduced compared to broadcast wireless channels. On the other hand, for the short-distance PA-user transmission, the near-field beam focusing can be employed to minimize the information leakage.
@@ V Numerical Case Studies
In this section, we present numerical examples to characterize the performance of PASS employing pinching beamforming and to compare them with other baselines.
@@ V-A Array Gains of PASS
The array gain of PASS is analyzed by considering a setup where multiple PAs, equally spaced along a single waveguide, are used to serve a single-antenna user; see [11, Figure 1]. This setup represents the most basic pinching beamforming scenario, and the corresponding results serve as a lower bound for more advanced designs.
@@ V-B Transmit Power Minimization by PASS
Although PASS has demonstrated the ability to achieve enhanced array gain, the design of the joint transmit and pinching beamforming for maximizing the PASS performance is a non-trivial task. The key challenges stem not only from the coupling between transmit and pinching beamforming but also from the combined impact of all PAs on both the amplitude and phase of the signals. To tackle these challenges, we propose leveraging penalty dual decomposition (PDD) to effectively decompose the optimization parameters.
To evaluate the effectiveness of PDD in enhancing the performance of multi-user PASS, we examine a transmit power minimization problem constrained by minimum signal-to-interference-plus-noise ratio (SINR) requirement. We assume a 15 GHz system where five 50-meter waveguides are deployed in parallel with a spacing of 5 meters, serving four users randomly distributed 15-20 meters away from the BS. Each waveguide has 6 PAs, and a sub-connected multiplexing transmission architecture is used. For comparison, we consider two baseline systems: fully-digital MIMO with 5 antennas and massive MIMO with 30 antennas and 5 RF chains. The numerical results obtained by the Monte Carlo method are given in Fig. 6, where the number of available PA positions for discrete activation is set to 200. These results highlight the significant performance gains of PASS in reducing transmit power, even when compared to massive MIMO with a more complex transmission architecture. Additionally, it is interesting to observe that the performance loss caused by the proportional power model is negligible compared to the equal power model.
@@ VI Conclusions and Open challenges
In this article, the promising concept of PASS was introduced, highlighting its great potential for enabling efficient wireless communications over the ‘last meter’. The basic physical principle, signal model, and communication design of PASS have been discussed. By leveraging the novel pinching beamforming, two practical transmission architectures were developed for PASS, each having representative advantages and limitations. Moreover, a range of promising applications for employing PASS in wireless networks were explored. Numerical case studies demonstrated the superiority of PASS over conventional multi-antenna transmission. As PASS remains in its early stages of research, there are numerous open challenges and opportunities to unlock its full potential. Several of these key research directions are exemplified below:
  • Spatial Modeling for PASS with Stochastic Geometry: Stochastic geometry provides a powerful framework for evaluating the average performance of wireless networks, which can be used to determine key design metrics such as the required number of PAs, the ratio between PAs and users, and optimal placement strategies. However, the unique features of PASS—such as near-field effects, LoS-dominated links, and location-flexible pinching—necessitate the development of more accurate spatial models to capture the spatial randomness, which can be a promising and interesting research direction.
  • Machine Learning Empowered PASS: Machine learning empowers both channel estimation and fast pinching for PASS. Owing to the large spatial extent, PASS may observe non-stationary multi-path characteristics in different serving areas. Generative learning can capture the spatial non-stationarity and reconstruct large-scale PASS channels. Furthermore, low-complexity learning-based designs are required for user-centric pinching of a large number of PAs, thus manipulating PAs for both path loss compensation and constructive interference.
  • Advanced Hardware Designs for PASS: Pinching beamforming plays a critical role in determining the flexibility of PASS and the maximum achievable performance gain. Advanced PASS hardware designs are essential in this context, posing several challenges, such as real-time and continuous control of PA movements/activation, adjustable power leakage of PAs, mitigation of unwanted power losses, and desired dielectric materials. This requires multi-disciplinary research efforts.

### v2

@@ IV-A PASS for Ubiquitous Indoor and Outdoor Coverage
The low-complexity structure of PASS enables it to be easily deployed on diverse structures in both outdoor and indoor scenarios, such as building facades, rooftops, roadside guardrails, and indoor ceilings or walls, as depicted in Fig. 4. PASS address the key challenge of conventional wireless communication, where communication links are often partially or totally blocked by obstacles and suffer from significant long-distance path loss due to those remotely deployed BSs. In particular, PASS establishes LoS-dominated and short-distance wireless channels with the widespread deployment of waveguides and the optimal activation of PAs near the terminals. By preventing wireless signals from being wasted or even interfering other users, PASS is capable of facilitating ubiquitous outdoor and indoor coverage in a cost-efficient and energy-efficient manner. Several attractive applications of PASS in indoor and outdoor scenarios will be discussed below.
On the one hand, PASS is quite appealing in indoor scenarios. As shown in the top right of Fig. 4, PASS installed on the ceiling can flexibly activate PAs to provide wireless communications for users in different rooms. This mitigates the severe penetration losses suffered by wireless signals transmitted by outdoor BSs or indoor access points (APs), particularly at mmWave and THz frequencies. Moreover, by exploiting the near-field channel between PAs and users, efficient beam focusing can be achieved. This not only enhances the desired signal strength but also reduces interference to other users. On the other hand, the widely distributed PASS can also be utilized in outdoor mobile scenarios, such as aerial/space communications and vehicular communications, as illustrated in the middle of Fig. 4. For example, recall the fact that conventional BSs with down-tilt antennas are typically inefficient for aerial communications. As a remedy, PASS installed on rooftops and facades can realize tailored upward beamforming for communications and sensing to effectively support medium- and high-altitude aircrafts as well as low-altitude unmanned aerial vehicles (UAVs) or other high-altitude platforms. As a further advancement, PASS can be employed in space communications. By integrating multiple waveguides and numerous PAs, PASS can form a large-scale antenna array to combat the severe path loss typically encountered in high-frequency satellite communications. Moreover, with PASS uniformly deployed along roadside guardrails, efficient vehicular communications can be achieved by adaptively activating PAs matching with vehicular movement. This is particularly beneficial for autonomous driving in future wireless networks.
@@ IV-B Interplay between PASS and NOMA
The combination of PASS and non-orthogonal multiple access (NOMA) provides a “win-win” solution for 6G and beyond. On the one hand, since multiple PAs on a specific waveguide can only be fed with the same set of signals, how to serve multiple users simultaneously becomes a critical challenge. To address this, it is natural to use NOMA to multiplex multiple users’ signal in the power domain over the same radio resource [9]. On the other hand, PASS also facilitates flexible NOMA operations. Recall the fact that NOMA relies on successive interference cancellation (SIC) to handle co-channel interference and therefore SIC order significantly affects the communication performance achieved by each NOMA user. Conventionally, a fixed distance-based SIC order is employed for NOMA, where users closer to the BS, so called strong users, are assigned higher SIC orders due to better channel conditions. In contrast, as shown in the bottom right of Fig. 4, PASS-based NOMA enables a more flexible SIC operation. This is achieved by dynamically adjusting users’ channel conditions through the allocation of PAs and the design of their activation locations. For example, users originally farther from the BS/AP can have higher channel conditions with more numbers of PAs allocated. By doing so, more customized wireless services can be provided by the PASS-based NOMA.
@@ IV-C PASS for NISE/ISAC
In addition to enhancing communication, PASS significantly improves sensing quality and coverage. Since the links between PAs and sensing targets follow the near-field model, PASS enables near-field sensing (NISE) to precisely estimate angular and range information, as well as target velocity [15]. As illustrated in the top left of Fig. 4, two practical structures can be employed: ISAC-PASS and NISE-PASS. (1) ISAC-PASS allocates part of the waveguides to transmit communication signals while using the rest to receive echo-sensing signals, enabling simultaneous near-field communication and sensing within the area of interest. (2) NISE-PASS, equipped exclusively with receiving waveguides, operates in a low-power mode to facilitate NISE. By leveraging the widespread deployment of PASS and successively or periodically activating sensing PAs, ubiquitous and highly accurate NISE can be achieved. This capability is particularly beneficial for real-world applications, such as tracking target trajectories or monitoring malicious entities.
@@ IV-D PASS for WPT
Wireless power transfer (WPT) is an efficient technology that uses RF signals to supply power to energy-constrained devices and extend their operational lifetime, especially for Internet-of-Things (IoT) and sensor networks. However, conventional WPT is hindered by the challenge of near-far unfairness, where devices located farther from the wireless power-supply BS or AP experience reduced energy-harvesting efficiency due to degraded channel conditions. As shown in the middle right of Fig. 4, this issue can be resolved through PASS-empowered WPT. In particular, tailored PAs can be activated right to each device, which ensures high-quality WPT for all devices, regardless of their locations.
@@ IV-E PASS for PLS
PASS can also enhance secure information transmission in the context of physical layer security (PLS). As shown in the bottom left of Fig. 4, PASS enables legitimate information transmission to effectively bypass the eavesdropper. On the one hand, since signal transmission predominantly occurs within the waveguide of PASS, the likelihood of legitimate information being eavesdropped is significantly reduced compared to broadcast wireless channels. On the other hand, for the short-distance PA-user transmission, the near-field beam focusing can be employed to minimize the information leakage.
@@ V Numerical Case Studies
In this section, we present numerical examples to characterize the performance of PASS employing pinching beamforming and to compare them with the conventional baselines.
@@ V-A Array Gains of PASS
The array gain of PASS is analyzed by considering a setup where multiple PAs, equally spaced along a single waveguide, are used to serve a single-antenna user; see [11, Figure 1]. This setup represents the most fundamental pinching beamforming scenario, and the corresponding results serve as a lower bound for those more advanced designs.
@@ V-B Transmit Power Minimization by PASS
Although PASS has demonstrated the ability to achieve enhanced array gains, the design of the joint transmit and pinching beamforming for maximizing the PASS performance is a non-trivial task. The key challenges stem not only from the coupling between transmit and pinching beamforming but also from the combined impact of all PAs on both the amplitude and phase of the signals. To tackle these challenges, we propose leveraging penalty dual decomposition (PDD) to effectively decompose the optimization parameters.
To evaluate the effectiveness of PDD in enhancing the performance of multi-user PASS, we examine a transmit power minimization problem constrained by the minimum signal-to-interference-plus-noise ratio (SINR) requirement. We assume a 15 GHz system where five 50-meter waveguides are deployed in parallel with a spacing of 5 meters, serving four users randomly distributed 15-20 meters away from the BS. Each waveguide has 6 PAs, and a sub-connected multiplexing transmission architecture is used. For comparison, we consider two baseline systems: fully-digital MIMO with 5 antennas and massive MIMO with 30 antennas and 5 RF chains. The numerical results obtained by the Monte Carlo method are illustrated in Fig. 6, where the number of available PA positions for discrete activation is set to 200. These results highlight the significant performance gains of PASS in reducing transmit power, even when compared to massive MIMO with a more complex transmission architecture. Additionally, it is interesting to observe that the performance loss caused by the proportional power model is negligible compared to the equal power model.
@@ VI Conclusions and Open challenges
In this article, the promising concept of PASS was introduced, highlighting their great potential for enabling efficient wireless communications over the “last meter”. The basic physical principle, signal model, and communication design of PASS have been discussed. By leveraging the novel pinching beamforming, two practical transmission architectures were developed for PASS, whose advantages and limitations were also described. Moreover, a range of promising applications for employing PASS in wireless networks were explored. Numerical case studies demonstrated the superiority of PASS over conventional multi-antenna transmission. As PASS remains in its early stages of research, there are numerous open challenges and opportunities to unlock its full potential. Several of these key research directions are exemplified below:
  • Spatial Modeling for PASS with Stochastic Geometry: Stochastic geometry provides a powerful framework for evaluating the average performance of wireless networks, which can be used to determine key design metrics, such as the required number of PAs, the ratio between PAs and users, and optimal placement strategies. However, the unique features of PASS—such as near-field effects, LoS-dominated links, and location-flexible pinching—necessitate the development of more accurate spatial models to capture the spatial randomness, which can be a promising and interesting research direction.
  • Machine Learning Empowered PASS: Machine learning empowers both channel estimation and fast pinching deployment for PASS. Owing to the large spatial extent, PASS may observe non-stationary multi-path characteristics in different serving areas. Generative learning can capture the spatial non-stationarity and reconstruct large-scale PASS channels. Furthermore, low-complexity learning-based designs are required for user-centric pinching of a large number of PAs, thus manipulating PAs for both path loss compensation and constructive interference.
  • Advanced Hardware Designs for PASS: Pinching beamforming plays a critical role in determining the flexibility of PASS and the maximum achievable performance gain. Advanced PASS hardware designs are essential in this context, posing several challenges, such as real-time and continuous control of PA movements or activation, active control of power leakage from PAs, mitigation of unwanted power losses, and desired dielectric materials. This requires multi-disciplinary research efforts.
Biographies
Yuanwei Liu (Fellow, IEEE) is a (tenured) Full Professor at The University of Hong Kong.
Zhaolin Wang (Member, IEEE) is a Post-Doctoral Researcher at Queen Mary University of London.
Xidong Mu (Member, IEEE) is a Lecturer with the Centre for Wireless Innovation (CWI), Queen’s University Belfast.
Chongjun Ouyang (Member, IEEE) is a Post-Doctoral Researcher at Queen Mary University of London.
Xiaoxia Xu (Member, IEEE) is a Post-Doctoral Researcher at Queen Mary University of London.
Zhiguo Ding (Fellow, IEEE) is a Professor in Communications at the University of Manchester and a Distinguished Adjunct Professor at Khalifa University.